# Privacy Amplification for Synthetic data using Range Restriction

### v1

@@ 2.1 Range-averaged Formal Privacy
Our definition of the probability $1-\lambda_{i}$ that unit $i$ is included in the sensitive range of the data space led us to deconstruct the likelihood into a sensitive and non-sensitive component. We used the likelihood decomposition to define a new mechanism under an implied new privacy standard that is amplified relative to Differential privacy because it focuses solely on the sensitive portion of the likelihood contribution (by conditioning on known information about sensitive data ranges).
We now make the definition of this new, amplified privacy standard explicit with a warm-up focus on the non-weighted posterior mechanism followed by a generalization to the risk-weighted pseudo posterior mechanism.
Our new privacy standard conditions the posterior mechanism on the set of record-indexed ranges or balls chosen by the owner of the closely-held data.
Let data $(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$, where each $x_{i}\in\mathbb{R}$. Suppose ranges, $R_{1},…,R_{n}\in\mathcal{X}^{n}\subset\mathbb{R}^{n}$ are designated to be sensitive by the owner of the closely-held data. The specification of this range gives rise to a probability that a datum will not be contained in the sensitive range, $\lambda(X,R)$, defined for $X\in\mathcal{X}$, $R\in\mathbb{R}$ and produces likelihood weight, $\alpha^{\ast}(X,R)$.
We then use the same definition of event probabilities as in DP and aDP, but now conditioning on $\mathbf{R}$.
@@
where $\bm{\lambda}\left(\mathbf{x}\right)=\left(\lambda_{1}(\mathbf{x}),\ldots,\lambda_{n}(\mathbf{x})\right)$ for $\lambda_{i}$ defined as the probability that datum $x_{i}$ lies outside the range deemed sensitive by the owner of the data in Equation 2. Define $\delta(\mathbf{x},\mathbf{x}^{{}^{\prime}})$ to be the number of records that differ between $\mathbf{x}$ and $\mathbf{x}^{{}^{\prime}}$ (with all other records being identical in both datasets).
@@
We don’t marginalize or maximize over $\mathbf{R}$, but condition on it because it is known information. $\lambda(X,R)$ marginalizes over data generating parameters $\theta\in\Theta$ that implicitly depends on model, $\mathcal{M}$. We suppress that conditioning because the entire set-up implicitly depends on $\mathcal{M}$.
We note that the mechanism under the range-averaged privacy standard is simply the posterior distribution. So, the appearance of $1-\lambda_{i}$ reflects the portion of the likelihood contribution that is sensitive and needs to be protected. We see this by examining the resulting sensitivity under range-averaged privacy,
$\Delta_{\bm{\lambda},\mathbf{x}}$ $=\max_{\theta\in\{\xi(\theta\mid\bm{x})\}_{m}}\max_{i\in 1,\ldots,n}|f_{\theta_{m}}(y_{i})-\lambda_{i}\times f_{\theta_{m}}(y_{i})|$ (10)
$=\max_{\theta\in\{\xi(\theta\mid\bm{x})\}_{m}}\max_{i\in 1,\ldots,n}|\left(1-\lambda_{i}\right)\times f_{\theta_{m}}(y_{i})|\leq\Delta_{\bm{\mathbf{x}}},$ (11)
where the correct sensitivity, $\Delta_{\bm{\lambda},\mathbf{x}}$, is achieved by subtracting the non-sensitive portion of the data range in Equation 10 from the full data log-likelihood.
@@
$\mathop{\sup}_{\mathbf{x}\in\mathcal{X}^{n},\mathbf{x}\in\mathcal{X}^{n-1}:\delta(\mathbf{x},\mathbf{x})=1}\mathop{\sup}_{B\in\beta_{\Theta}}\frac{\xi^{\bm{\lambda}^{c}\bm{\alpha}(\mathbf{x})}(B\mid\mathbf{x},\mathbf{R})}{\xi^{\bm{\lambda}^{c}\bm{\alpha}(\mathbf{x})}(B\mid\mathbf{x},\mathbf{R})}\leq e^{\epsilon},$
The mechanism associated with the range restricted pseudo posterior privacy standard is the pseudo posterior mechanism with the likelihood contribution for each unit $i\in(1,\ldots,n)$ exponentiated by $\alpha^{\ast}_{i}$.
This new range-averaged privacy standard can be viewed as relaxation of Differential privacy because it averages over datasets to compute $\lambda_{i}$ (rather than bounds a worst case), but it is also an amplification because it conditions on non-sensitive ranges. Yet, beyond comparing to Differential privacy, the range-averaged privacy represents a coherent, formal standard in its own right (e.g., like the Pufferfish formal privacy standard of Kifer and Machanavajjhala (2014)). Both the range-averaged and Pufferfish privacy standards provide amplifcations of the privacy guarantee provided by Differential privacy by defining their standards to condition on additional information/assumptions.
One may gain insight into the relative strength of the range-averaged privacy guarantee by focusing on the corner case of a highly risky record. Under the usual aDP guarantee (with no range restriction) this datum’s likelihood contribution would be assigned a low value for the risk weight, $\alpha$. If, however, if the datum for this record would rarely fall within the sensitive range under the generating distribution for this record then the likelihood weight, $\alpha^{\ast}>\alpha$, with no loss of privacy protection since that otherwise risky record is deemed sensitive with a low probability, $1-\lambda$.
@@
$\forall\mathbf{x}\in\mathcal{X}^{n},\mathbf{x}\in\mathcal{X}^{n-1}:\delta(\mathbf{x},\mathbf{x})=1,B\in\beta_{\Theta}$ (where $\beta_{\Theta}$ is the $\sigma$-algebra of measurable sets on $\Theta$) under $\bm{\alpha}(\cdot)$ with $\Delta_{\bm{\alpha},\bm{\lambda},\mathbf{x}}>0$,
$\mathop{\sup}_{B\in\beta_{\Theta}}\,\,\,\frac{\xi^{\bm{\lambda}^{c}\bm{\alpha}(\mathbf{x})}(B\mid\mathbf{x})}{\xi^{\bm{\lambda}^{c}\bm{\alpha}(\mathbf{x})}(B\mid\mathbf{x})}\leq\exp(2\Delta_{\bm{\alpha},\bm{\lambda},\mathbf{x}}),$ (12)
i.e. the pseudo posterior $\xi^{\bm{\lambda}^{c}\bm{\alpha}(\mathbf{x})}(\cdot\mid\mathbf{x})$ has local privacy guarantee $2\Delta_{\bm{\alpha},\bm{\lambda},\mathbf{x}}$.
@@
We note that the guarantee of Theorem 2 is local to dataset $\mathbf{x}\in\mathcal{X}^{n}$. The global sensitivity, $\Delta_{\bm{\alpha},\bm{\lambda}}=\mathop{\sup_{\mathbf{x}\in\mathcal{X}^{n}}}\Delta_{\bm{\alpha},\bm{\lambda},\mathbf{x}}$ is constructed as the upper bound of the sensitivities for all datasets $\mathbf{x}\in\mathcal{X}^{n}$. As reviewed in Section 1, Savitsky et al. (2022) demonstrate that for any realized $\mathbf{x}$ under a data generating process that $\Delta_{\bm{\alpha},\bm{\lambda},\mathbf{x}}$ contracts onto $\Delta_{\bm{\alpha},\bm{\lambda}}$ for sample size, $n$, sufficiently large (at $\mathcal{O}(n^{-\frac{1}{2}})$). They label their relaxed privacy standard as asymptotic Differential privacy (aDP). We may extend the large sample aDP bound, $\epsilon\leq 2\Delta_{\bm{\alpha}}$, to the pseudo posterior mechanism under our amplified range-averaged privacy standard to achieve a new $\epsilon\leq 2\Delta_{\bm{\alpha},\bm{\lambda}}$ range-averaged asymptotic Differential privacy (raaDP) as the limiting point of $\Delta_{\bm{\alpha},\bm{\lambda},\mathbf{x}}$ given in Equation 7.
@@ 3 Truncation of Record Inclusion in Sensitive Range
The range-averaged privacy standard utilizes what may be viewed as the “maximum” information about the sensitive range because it both uses the range, itself, and also the frequency with which the data values for a record appear in that sensitive range over repeated sampling from the data generating distribution. Even if the sensitive range itself is relatively wide, we are able to induce relatively less distortion to achieve a targeted privacy guarantee if the probability $1-\lambda_{i}$ that generated datum values for record $i$ will lie inside that range is small.
The range-averaged privacy standard produces a probability that the generating distribution of each datum $i$ will lie inside the sensitive range using a model, $\mathcal{M}$, estimated on the closely-held data, rather than declaring each unit as lying in the sensitive range or not. So, we allocate a $1-\lambda_{i}$ proportion of the likelihood for record $i$ as sensitive.
Perhaps a more conservative approach would be to focus on using the “minimum” information about the sensitive range. We declare the datum for a record as either lying inside or outside of that range and condition our use of the model distribution only on endpoints of the sensitive interval (rather than the distribution of data values within the sensitive range).
@@
Since the second pseudo likelihood component on the right hand most term is viewed as acceptable to expose to the public, we only use the first component (on the left hand term) to assess privacy risk.
We develop an alternative privacy standard by declaring a record’s datum value to lie inside or outside of a sensitive data range (with probability $1$), and then decompose the (pseudo likelihood) into two components based on interval censoring and its complement, interval truncation.
@@
$p^{I}(x_{i}|\theta,a,b)=\left\{\begin{array}[]{rl}P(b\times x_{i}|\theta)-P(a\times x_{i}|\theta),&x_{i}\in[a\times x_{i},b\times x_{i}]\\ p(x_{i}|\theta),&x\notin[a\times x_{i},b\times x_{i}]\end{array}\right$.
with $P(\cdot|\theta)$ being the cdf under our model for the closely-held data. This likelihood represents information generally known because when $x_{i}$ is in the sensitive range the edges of that range are known. By contrast, when $x_{i}$ is not in the sensitive range there is no need to privatize.
The sensitive complement (i.e. truncated) likelihood term is specified by:
$p^{I^{c}}(x_{i}|\theta,a,b)=\left\{\begin{array}[]{rl}p(x_{i}|\theta)/\left(P(b\times x_{i}|\theta)-P(a\times x_{i}|\theta)\right),&x_{i}\in[a\times x_{i},b\times x_{i}]\\ 1,&x_{i}\notin[a\times x_{i},b\times x_{i}]\end{array}\right$.
We define both the sensitive and non-sensitive components of the likelihood by allocating datum $x_{i}$ dichotomously to the sensitive region. If a record is assigned to the sensitive region of the likelihood complement, then we would fully apply our pseudo posterior risk-based weight, $\alpha_{i}$, to construct our pseudo posterior mechanism. The known information is not conveyed through adjusting the privacy weight, $\alpha_{i}$, in the the likelihood as is done in the range-averaging mechanism, but rather dividing the likelihood, $p(x_{i}\mid\theta)$, by the difference in the cdf values at the end points, $\left(P(b\times x_{i}|\theta)-P(a\times x_{i}|\theta)\right)$. Our resulting pseudo posterior mechanism is constructed as:
$p^{I^{c},\alpha_{i}}(x_{i}|\theta,a,b)=\left\{\begin{array}[]{rl}p(x_{i}|\theta)^{\alpha_{i}}/\left(P(b\times x_{i}|\theta)-P(a\times x_{i}|\theta)\right),&x\in[a\times x_{i},b\times x_{i}]\\ 1,&x_{i}\notin[a\times x_{i},b\times x_{i}]\end{array}\right$.
@@
$p_{\bm{\theta}}^{\bm{\alpha}}(\mathbf{x})=p_{\bm{\theta}}^{\bm{I}^{c}\bm{\alpha}}(\mathbf{x})p_{\bm{\theta}}^{\bm{I}}(\mathbf{x})=\mathop{\prod}_{i=1}^{n}p^{I^{c},\alpha_{i}}(x_{i}|\theta,a,b)\mathop{\prod}_{i=1}^{n}p^{I}(x_{i}|\theta,a,b).$ (14)
In practice, one will construct their sensitive range to be centered on or surrounding the datum value, $x_{i}\in[a\times x_{i},b\times x_{i}]$ so that $x_{i}$ always lies in the sensitive range. Then our sensitivity calculation focuses on the first line of the complement:
@@
The non-sensitive portion of the likelihood is expressed under range-averaged privacy through exponentiating the datum likelihood contribution by $\lambda_{i}$. The larger is $0\leq\lambda_{i}\leq 1$ the greater the portion of the datum likelihood admitted to the pseudo posterior model. We may interpret probability $\lambda_{i}$ from range-average privacy as the effective sample size or fraction of information the data producer thinks is “public” and thus does not need to be protected. By contrast, one divides the likelihood by the $P(b\times x_{i}|\theta)-P(a\times x_{i}|\theta)$ under the range-truncation privacy standard. For any finite $a$ and $b$, the difference in cdf functions is <1<1, which magnifies the likelihood as compared to an unrestricted range (in which case the divisor is exactly $1$). One may think of relatively risky records as located in the tails that expresses small likelihood values and the divisor would increase those values, which partially mitigates the $\alpha$ down-weighting. We expect $\Delta_{\bm{\alpha},\bm{I},\mathbf{x}}\leq\Delta_{\bm{\alpha},\mathbf{x}}$. When the interval $a\times x_{i},b\times x_{i}]$ widens into the whole range, the second term is $\log(1)=0$, which causes the range-truncated privacy standard to revert to the aDP privacy standard.

### v2

@@ 2.1 Range-averaged formal privacy
Our definition of the probability $(1-\lambda_{i})$ that unit $i$ is included in the sensitive range of the data space leads us to deconstruct the likelihood into a sensitive and non-sensitive component. We use the likelihood decomposition to define a new mechanism under an implied new privacy standard that is amplified relative to DP because it focuses solely on the sensitive portion of the likelihood contribution (by conditioning on known information about sensitive data ranges).
We now make the definition of this new, amplified privacy standard explicit with a warm-up focus on the non-weighted posterior mechanism followed by a generalization to the risk-weighted pseudo posterior mechanism. Our new privacy standard conditions the posterior mechanism on the set of record-indexed ranges or balls chosen by the owner of the closely-held data.
Let data $(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$, where each $x_{i}\in\mathbb{R}$. Suppose ranges, $R_{1},…,R_{n}\in\mathcal{X}^{n}\subset\mathbb{R}^{n}$ are designated to be sensitive by the owner of the closely-held data. The specification of this range gives rise to a probability that a datum will not be contained in the sensitive range, $\lambda(X,R)$, defined for $X\in\mathcal{X}$, $R\in\mathbb{R}$ and produces likelihood weight, $\alpha^{\ast}(X,R)$. We then use the same definition of event probabilities as in DP and aDP, but now conditioning on $\mathbf{R}$.
@@
where $\bm{\lambda}\left(\mathbf{x}\right)=\left(\lambda_{1}(\mathbf{x}),\ldots,\lambda_{n}(\mathbf{x})\right)$ for $\lambda_{i}$ defined as the probability that datum $x_{i}$ lies outside the range deemed sensitive by the owner of the data in Equation (3). Define $\delta(\mathbf{x},\mathbf{x}^{{}^{\prime}})$ to be the number of records that differ between $\mathbf{x}$ and $\mathbf{x}^{{}^{\prime}}$ (with all other records being identical in both datasets).
@@
We do not marginalize or maximize over $\mathbf{R}$, but condition on it because it is known information. $\lambda(X,R)$ marginalizes over data generating parameters $\theta\in\Theta$ that implicitly depends on model, $\mathcal{M}$. We suppress that conditioning because the entire set-up implicitly depends on $\mathcal{M}$.
We note that the mechanism under the range-averaged privacy standard is simply the posterior distribution. Therefore, the appearance of $(1-\lambda_{i})$ reflects the portion of the likelihood contribution that is sensitive and needs to be protected. We see this by examining the resulting sensitivity under range-averaged privacy,
$\Delta_{\bm{\lambda},\mathbf{x}}$ $=\max_{\theta\in\{\xi(\theta\mid\bm{x})\}_{m}}\max_{i\in 1,\ldots,n}|f_{\theta_{m}}(x_{i})-\lambda_{i}\times f_{\theta_{m}}(x_{i})|$ (11)
$=\max_{\theta\in\{\xi(\theta\mid\bm{x})\}_{m}}\max_{i\in 1,\ldots,n}|\left(1-\lambda_{i}\right)\times f_{\theta_{m}}(x_{i})|\leq\Delta_{\bm{\mathbf{x}}},$ (12)
where the correct sensitivity, $\Delta_{\bm{\lambda},\mathbf{x}}$, is achieved by subtracting the non-sensitive portion of the data range in Equation (11) from the full data log-likelihood.
@@
$\mathop{\sup}_{\mathbf{x}\in\mathcal{X}^{n},\mathbf{x}^{\prime}\in\mathcal{X}^{n-1}:\delta(\mathbf{x},\mathbf{x}^{\prime})=1}\mathop{\sup}_{B\in\beta_{\Theta}}\frac{\xi^{\bm{\lambda}^{c}\bm{\alpha}(\mathbf{x})}(B\mid\mathbf{x},\mathbf{R})}{\xi^{\bm{\lambda}^{c}\bm{\alpha}(\mathbf{x}^{\prime})}(B\mid\mathbf{x}^{\prime},\mathbf{R})}\leq e^{\epsilon},$
The mechanism associated with the range-restricted pseudo posterior privacy standard is the pseudo posterior mechanism with the likelihood contribution for each unit $i\in(1,\ldots,n)$ exponentiated by $\alpha^{\ast}_{i}$.
This new range-averaged privacy standard can be viewed as relaxation of DP because it averages over datasets to compute $\lambda_{i}$ (rather than bounds a worst case), but it is also an amplification because it conditions on non-sensitive ranges. Yet, beyond comparing to DP, the range-averaged privacy represents a coherent, formal standard in its own right (e.g., like the Pufferfish formal privacy standard of Kifer and Machanavajjhala (2014)). Both the range-averaged and Pufferfish privacy standards provide amplifications of the privacy guarantee provided by DP by defining their standards to condition on additional information/assumptions.
One may gain insight into the relative strength of the range-averaged privacy guarantee by focusing on the corner case of a highly risky record. Under the usual aDP guarantee (with no range restriction) this datum’s likelihood contribution would be assigned a low value for the risk weight, $\alpha$. However, if the datum for this record would rarely fall within the sensitive range under the generating distribution for this record then the likelihood weight, $\alpha^{\ast}_{i}>\alpha_{i}$, with no loss of privacy protection since that otherwise risky record is deemed sensitive with a low probability, $(1-\lambda_{i})$.
@@
$\forall\mathbf{x}\in\mathcal{X}^{n},\mathbf{x}^{\prime}\in\mathcal{X}^{n-1}:\delta(\mathbf{x},\mathbf{x}^{\prime})=1,B\in\beta_{\Theta}$ (where $\beta_{\Theta}$ is the $\sigma$-algebra of measurable sets on $\Theta$) under $\bm{\alpha}(\cdot)$ with $\Delta_{\bm{\alpha},\bm{\lambda},\mathbf{x}}>0$,
$\mathop{\sup}_{B\in\beta_{\Theta}}\,\,\,\frac{\xi^{\bm{\lambda}^{c}\bm{\alpha}(\mathbf{x})}(B\mid\mathbf{x})}{\xi^{\bm{\lambda}^{c}\bm{\alpha}(\mathbf{x}^{\prime})}(B\mid\mathbf{x}^{\prime})}\leq\exp(2\Delta_{\bm{\alpha},\bm{\lambda},\mathbf{x}}),$ (13)
i.e., the pseudo posterior $\xi^{\bm{\lambda}^{c}\bm{\alpha}(\mathbf{x})}(\cdot\mid\mathbf{x})$ has local privacy guarantee $2\Delta_{\bm{\alpha},\bm{\lambda},\mathbf{x}}$.
@@
We note that the guarantee of Theorem 2 is local to dataset $\mathbf{x}\in\mathcal{X}^{n}$. The global sensitivity, $\Delta_{\bm{\alpha},\bm{\lambda}}=\mathop{\sup_{\mathbf{x}\in\mathcal{X}^{n}}}\Delta_{\bm{\alpha},\bm{\lambda},\mathbf{x}}$ is constructed as the upper bound of the sensitivities for all datasets $\mathbf{x}\in\mathcal{X}^{n}$. As reviewed in Section 1.1, Savitsky et al. (2022) demonstrate that for any realized $\mathbf{x}$ under a data generating process that $\Delta_{\bm{\alpha},\bm{\lambda},\mathbf{x}}$ contracts onto $\Delta_{\bm{\alpha},\bm{\lambda}}$ for sample size, $n$, sufficiently large (at $\mathcal{O}(n^{-\frac{1}{2}})$). This relaxed privacy standard is hence labeled as aDP where a stands for asymptotic. We may extend the large sample aDP bound, $\epsilon\leq 2\Delta_{\bm{\alpha}}$, to the pseudo posterior mechanism under our amplified range-averaged privacy standard to achieve a new $\epsilon\leq 2\Delta_{\bm{\alpha},\bm{\lambda}}$ range-averaged aDP as the limiting point of $\Delta_{\bm{\alpha},\bm{\lambda},\mathbf{x}}$ given in Equation (8).
@@ 3 Truncation of record inclusion in sensitive range
The range-averaged privacy standard introduced in Section 2 utilizes what may be viewed as the “maximum” information about the sensitive range because it both uses the range and the frequency with which the data values for a record appear in that sensitive range over repeated sampling from the data generating distribution. Even if the sensitive range itself is relatively wide, we are able to induce relatively less distortion to achieve a targeted privacy guarantee if the probability $(1-\lambda_{i})$ that generates datum values for record $i$ will lie inside that range is small. The range-averaged privacy standard produces a probability that the generating distribution of each datum $i$ will lie inside the sensitive range using a model, $\mathcal{M}$, estimated on the closely-held data, rather than declaring each unit as lying in the sensitive range or not. Therefore, we allocate a $(1-\lambda_{i})$ proportion of the likelihood for record $i$ as sensitive.
Perhaps a more conservative approach would be to focus on using the “minimum” information about the sensitive range. We declare the datum for a record as either lying inside or outside of that range and condition our use of the model distribution only on endpoints of the sensitive interval, rather than the distribution of data values within the sensitive range as in the range-averaged standard.
@@
Since the second pseudo likelihood component on the right hand most term in Equation (14) is viewed as acceptable to expose to the public, we only use the first component to assess privacy risk.
We develop an alternative privacy standard by declaring a record’s datum value to lie inside or outside of a sensitive data range with probability $1$, and then decompose the pseudo likelihood into two components based on interval censoring and its complement, interval truncation.
@@
$p^{I}(x_{i}|\theta,a,b)=\left\{\begin{array}[]{rl}P(b\times x_{i}\mid\theta)-P(a\times x_{i}\mid\theta),&x_{i}\in[a\times x_{i},b\times x_{i}]\\ p(x_{i}\mid\theta),&x_{i}\notin[a\times x_{i},b\times x_{i}],\end{array}\right$.
with $P(\cdot\mid\theta)$ being the cumulative density function (CDF) under our model for the closely-held data. This likelihood represents information generally known because when $x_{i}$ is in the sensitive range and the endpoints of that range are known. By contrast, when $x_{i}$ is not in the sensitive range there is no need to provide privacy protection.
The sensitive complement (i.e., truncated) likelihood term is specified by
$p^{I^{c}}(x_{i}|\theta,a,b)=\left\{\begin{array}[]{rl}p(x_{i}|\theta)/\left(P(b\times x_{i}|\theta)-P(a\times x_{i}|\theta)\right),&x_{i}\in[a\times x_{i},b\times x_{i}]\\ 1,&x_{i}\notin[a\times x_{i},b\times x_{i}].\end{array}\right$.
We define both the sensitive and non-sensitive components of the likelihood by allocating datum $x_{i}$ dichotomously to the sensitive region. If a record is assigned to the sensitive region of the likelihood complement, then we would fully apply our pseudo posterior risk-based weight, $\alpha_{i}$, to construct our pseudo posterior mechanism. The known information is not conveyed through adjusting the privacy weight, $\alpha_{i}$, in the the likelihood as is done in the range-averaging standard, but rather dividing the likelihood, $p(x_{i}\mid\theta)$, by the difference in the CDF values at the endpoints, $\left(P(b\times x_{i}|\theta)-P(a\times x_{i}|\theta)\right)$. Our resulting pseudo posterior mechanism is constructed as
$p^{I^{c},\alpha_{i}}(x_{i}|\theta,a,b)=\left\{\begin{array}[]{rl}p(x_{i}|\theta)^{\alpha_{i}}/\left(P(b\times x_{i}|\theta)-P(a\times x_{i}|\theta)\right),&x_{i}\in[a\times x_{i},b\times x_{i}]\\ 1,&x_{i}\notin[a\times x_{i},b\times x_{i}]\end{array}\right$.
@@
$p_{\bm{\theta}}^{\bm{\alpha}}(\mathbf{x})=p_{\bm{\theta}}^{\bm{I}^{c}\bm{\alpha}}(\mathbf{x})p_{\bm{\theta}}^{\bm{I}}(\mathbf{x})=\mathop{\prod}_{i=1}^{n}p^{I^{c},\alpha_{i}}(x_{i}\mid\theta,a,b)\mathop{\prod}_{i=1}^{n}p^{I}(x_{i}\mid\theta,a,b).$ (15)
In practice, one will construct their sensitive range to be centered on or surrounding the datum value, $x_{i}\in[a\times x_{i},b\times x_{i}]$ so that $x_{i}$ always lies in the sensitive range. Then our sensitivity calculation focuses on the first complement:
@@
Recall the non-sensitive portion of the likelihood is expressed under range-averaged privacy in Section 2 through exponentiating the datum likelihood contribution by $\lambda_{i}$. The larger is $0\leq\lambda_{i}\leq 1$ the greater the portion of the datum likelihood admitted to the pseudo posterior model. We may interpret probability $\lambda_{i}$ from range-average privacy as the effective sample size or fraction of information the data producer thinks is “public” and thus does not need to be protected. By contrast, one divides the likelihood by the $P(b\times x_{i}\mid\theta)-P(a\times x_{i}\mid\theta)$ under the range-truncation privacy standard. For any finite $a$ and $b$, the difference in CDF functions is <1<1, which magnifies the likelihood as compared to an unrestricted range (in which case the divisor is exactly $1$). One may think of relatively risky records as located in the tails that expresses small likelihood values and the divisor would increase those values, which partially mitigates the $\bm{\alpha}$ down-weighting. We expect $\Delta_{\bm{\alpha},\bm{I},\mathbf{x}}\leq\Delta_{\bm{\alpha},\mathbf{x}}$. When the interval $a\times x_{i},b\times x_{i}]$ widens into the whole range, the second term is $\log(1)=0$, which causes the range-truncated privacy standard to revert to the aDP privacy standard.